OpenStudy (anonymous):

Find approximate solution for the equation in the interval 0 < x < pi . Rounded to three decimal places. tanx = 5 sinx

OpenStudy (tkhunny):

Have you considered factoring?

OpenStudy (anonymous):

how would i factor that?

OpenStudy (tkhunny):

\(\tan(x) - 5\sin(x) = 0\) \(\dfrac{\sin(x)}{\cos(x)} - 5\sin(x) = 0\) \(\sin(x)\cdot \left(\dfrac{1}{\cos(x)} - 5\right) = 0\)

OpenStudy (anonymous):

i dont understand

OpenStudy (saifoo.khan):

Which part? @essalytee

OpenStudy (anonymous):

i just dont understand the whole problem

OpenStudy (saifoo.khan):

Do you understand his step1?

OpenStudy (anonymous):

yes, i understand his steps but i dont know how to work out the problem after that

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

now after step 3, set sin(x) = 0 AND (1/cos x - 5) = 0 Separately.

OpenStudy (anonymous):

so the answer would be 0, 1.571?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

sin(x) = 0 will give x = 0 For the next one: cos(x) = 1/5 x = 1.37

OpenStudy (anonymous):

okay thank you! i did some wrong calculations for the second one!

OpenStudy (saifoo.khan):

No problem.